Keep 5 decimal places for intermediate steps and report your final answers in 2 decimal places (16 points)

Number of participants enrolled at baseline

Total Person Months

Number of new DVT cases by 24 months

OCP Use

Yes

500

6600

90

No

300

3800

40

Total

800

10400

130

  1. Calculate the risk ratio for DVT among exposed and unexposed women. (2 points)
Answer #1

Hey, welcome to Chegg. I hope my answer will be of some help to you. Thank you!

The probability of DVT in exposed (OCP using) women = Number of DVT cases/ Number of exposed participants

= 90/ 6600 = 0.01363

The probability of DVT in unexposed ( Non-OCP using) women = Number of DVT cases/ Number of unexposed participants

= 40/ 3800 = 0.01052

Risk Ratio can calculated by the ratio of probability of DVT in exposed participants to the probability of DVT in unexposed participants.

RR = (0.01363) ÷ (0.01052) = 1.29